# SupSup Ref

### v1

@@ Abstract
A long-standing goal of 3D human reconstruction is to create lifelike and fully detailed 3D humans from single images. The main challenge lies in inferring unknown human shapes, clothing, and texture information in areas not visible in the images. To address this, we propose SiTH, a novel pipeline that uniquely integrates an image-conditioned diffusion model into a 3D mesh reconstruction workflow. At the core of our method lies the decomposition of the ill-posed single-view reconstruction problem into hallucination and reconstruction subproblems. For the former, we employ a powerful generative diffusion model to hallucinate back appearances from the input images. For the latter, we leverage skinned body meshes as guidance to recover full-body texture meshes from the input and back-view images. Our designs enable training of the pipeline with only about 500 3D human scans while maintaining its generality and robustness. Extensive experiments and user studies on two 3D reconstruction benchmarks demonstrated the efficacy of our method in generating realistic, fully textured 3D humans from a diverse range of unseen images.
Figure 1: Single-view textured human reconstruction. We propose SiTH, a novel approach for creating high-quality and fully textured 3D human meshes from single images. We first hallucinate back-view appearances through an image-conditioned diffusion model, followed by the reconstruction of full-body textured meshes using both the front and back-view images. Our pipeline enables the creation of lifelike and diverse 3D humans from unseen photos (left) and AI-generated images (right). Best viewed in color.
@@ 1 Introduction
Reconstructing a fully textured human mesh from a single-view image presents an ill-posed problem with two major challenges. Firstly, the appearance information required for generating texture in unobserved regions is missing. Secondly, 3D information for mesh reconstruction, such as depth, surface, and body pose, becomes ambiguous in a 2D image. Previous efforts [44, 60, 2] attempted to tackle these challenges in a data-driven manner, focusing on generating image-mesh pairs to train neural networks. However, these approaches struggle with images featuring unseen appearances or poses (see Fig. 5), due to limited 3D human training data. More recent studies [45, 60, 54, 55] introduced additional 3D reasoning modules to enhance robustness against unseen poses. Yet, generating realistic and full-body textures from unseen appearances still remains an unsolved problem.
@@
To address the above challenges, we propose SiTH, a novel pipeline that uniquely integrates an image-condition diffusion model to reconstruct lifelike 3D textured humans from monocular images. At the core of our approach is the decomposition of the challenging single-view problem into two subproblems: generative back-view hallucination and mesh reconstruction. This decomposition enables us to exploit the generative capability of pretrained diffusion models to guide full-body mesh and texture reconstruction. The workflow is depicted in Fig. 1. Given a front-view image, the first stage involves hallucinating a perceptually consistent back-view image using image-conditioned diffusion. The second stage reconstructs full-body mesh and texture, utilizing both the front and back-view images as guidance.
@@
More specifically, we employ the generative capabilities of pretrained diffusion models (e.g. Stable Diffusion [41]) to infer unobserved back-view appearances for full-body 3D reconstruction. The primary challenge in ensuring the realism of 3D meshes lies in generating images that depict spatially aligned body shapes and perceptually consistent appearances with the input images. While diffusion models demonstrate impressive generative abilities with text conditioning, they are limited in producing desired back-view images using the frontal images as image conditions. To overcome this, we adapt the network architecture to enable conditioning on frontal images and introduce additional trainable components following ControlNet [58] to provide pose and mask control. To fully tailor this model to our task while retaining its original generative power, we apply adapter-based fine-tuning [15, 58] using multi-view images rendered from 3D human scans. Complementing this generative model, we develop a mesh reconstruction module to recover full-body textured mesh from front and back-view images. We follow prior work in handling 3D ambiguity through normal [45] and skinned body [54, 60] guidance. Compared to existing end-to-end methods, our two-stage pipeline can recover full-body textured meshes, including back-view details, and demonstrates robustness to unseen images (Fig. 1). It is noteworthy that the models for both subproblems are trained using the same public THuman2.0 [56] dataset, which consists of about 500 scans.
Furthermore, we create a single-view human reconstruction benchmark using a high-quality 3D human dataset, CustomHumans [14]. In this benchmark, we conducted detailed evaluations, ablations, and user studies, to compare our method with state-of-the-art approaches. Additionally, we demonstrated an application combining text-guided diffusion models, showing our potential for supporting 3D human creation. Our contributions are three-fold:
  • We propose SiTH, a single-view human reconstruction pipeline that generates lifelike and fully textured 3D humans using image-condition diffusion models.
  • By decomposing the complex, ill-posed problem, SiTH is robust to unseen images and can be effectively trained using publicly available 3D human scans.
  • SiTH can be seamlessly integrated with powerful text-to-image diffusion models for 3D human creation.
@@ 2 Related Work
We review methods that recover 3D human shapes, garments, and textures. One line of work utilizes neural fields [53] to represent clothed 3D humans. As a seminal work, Saito et al. [44] first proposed a data-driven method with pixel-aligned features. Its follow-up work PIFuHD [45] further improved this framework with high-res normal guidance. Follow-up approaches extend this framework with additional human body priors. For instance, PaMIR [60] and ICON [54] utilize skinned body models [28, 35] to guide 3D reconstruction. ARCH [16], ARCH++ [13], and CAR [25] transform global coordinates into the canonical coordinates to allow for reposing. PHOHRUM [2] and S3F [7] further disentangle shading and albedo to enable relighting. Another line of work replaces the neural representations with conventional Poisson surface reconstruction [20, 19]. ECON [55] and 2K2K [10] train normal and depth predictors to generate front and back 3D point clouds. The human mesh is obtained by combining these point clouds with body priors and 3D heuristics. However, the above models cannot produce realistic full-body texture and geometry in the unobserved regions. Our pipeline addresses this problem by incorporating a generative diffusion model into single-view human reconstruction.
@@
Diffusion models [41, 43, 40, 36] trained with large collections of images have demonstrated unprecedented capability in creating 3D objects with text prompts. Most prior work [37, 51, 26, 32, 11, 6] followed an optimization workflow to update 3D representations (e.g. NeRF [33], SDF tetrahedron [46]) via neural rendering [50] and a score distillation sampling (SDS) [37] loss. While some methods [57, 5, 17] applied this workflow to human bodies, only common subjects can be generated due to the ambiguity of text-conditioning. More recent work [27, 38, 48, 47] also tried to extend this workflow with more accurate image-conditioning. However, 3D humans require geometric details on clothing and facial expressions. In our experiments, we show that these methods struggle to recover such details and take a long optimization time. Most related to our work is Chupa [21], which also decomposes the pipeline into two stages. Note that Chupa does not generate texture and only uses texts for conditioning. Our methods address these issues by leveraging an image-conditioned diffusion model to create fully textured human meshes. More importantly, our method does not require a long optimization process.
@@
Foundation models [8, 3, 12, 22] trained on large-scale datasets have been shown to be adaptable to various downstream tasks. Following this trend, pretrained diffusion models [41, 43, 40, 36] have become common backbones for generative modeling. For instance, they can be customized by finetuning with a small collection of images [15, 23, 42]. ControlNet [58] introduced additional trainable plugins to enable image conditioning such as body skeletons. While these strategies have been widely adopted, none of them directly fit our objective. More relevant to our task is DreamPose [18], which utilizes DensePose [9] images as conditions to repose input images. However, it cannot handle out-of-distribution images due to overfitting. Zero-1-to-3 [27] finetunes a diffusion model with multi-view images to allow for viewpoint control. However, we show that viewpoint conditioning is not sufficient for generating consistent human bodies. Our model addresses this issue by providing accurate body pose and mask conditions for back-view hallucination.
@@ 3.1 Back-view Hallucination
Simply applying the LDM architecture to our task is not sufficient since our goal is to learn a conditional distribution of back-view images given an input conditional image. To this end, we make several adaptations to allow for image-conditioning as shown in Fig. 3. First, we utilize the pretrained CLIP [39] image encoder to extract image features from the front-view image (i.e., $I^{F}$). These image features are used for conditioning the LDM via cross-attention transformer layers, ensuring the output image shares a consistent appearance with the input image. Second, we follow the idea of ControlNet [58] and propose to use a UV map ($I^{B}_{UV}\in\mathbb{R}^{H\times W\times 3}$) and a silhouette mask ($I^{B}_{M}\in\mathbb{R}^{H\times W}$) from the back view as additional conditions. These conditional signals provide additional information that ensures the output image has a similar body shape and pose to the conditional input image.
@@
Another challenge in training an image-conditioned LDM is data. Training the model from scratch is infeasible due to the requirement of a large number of paired images rendered from 3D textured human scans. Inspired by the concept of learning from large-scale pretraining [8, 12], we build our image-conditioned LDM on top of a pretrained diffusion U-Net [41]. We follow the training strategy in [23, 58] to optimize cross-attention layers and ControlNet parameters while keeping most of the other parameters frozen (see Fig. 3). The design and training strategy of our diffusion model enables hallucinating plausible back-view images that are cosistent with the frontal inputs.
@@ 3.2 Human Mesh Reconstruction
$\mathbf{f}_{r,x}=\mathcal{B}(\mathbf{f}_{r},\pi(\mathbf{x}))=\mathcal{B}(G_{r}(I),\pi(\mathbf{x})),$ (5)
@@
$\mathbf{r}_{x}=H_{r}(\mathbf{f}_{r,x}^{F},\mathbf{f}_{r,x}^{B},\mathbf{p}).$ (8)
CAPE [30] CustomHumans [14]
@@
PIFu* [44] 3.350 0.755 29.283 0.9898 0.0198 2.939 0.780 31.911 0.9675 0.0454
PIFuHD* [45] 3.185 0.755 32.157 - - 2.651 0.785 36.311 - -
PaMIR [60] 3.405 0.742 25.475 0.9881 0.0140 3.591 0.753 25.716 0.9640 0.0348
ICON [54] 3.240 0.757 25.898 - - 3.190 0.763 26.857 - -
ECON [55] 3.084 0.765 26.960 - - 3.315 0.757 27.075 - -
SiTH (Ours) 2.530 0.790 31.256 0.9926 0.0111 2.443 0.805 33.061 0.9732 0.0291
Table 1: Single-view human reconstruction benchmarks. We report Chamfer distance (CD), normal consistency (NC), and f-score between ground truth and predicted meshes. To evaluate texture reconstruction quality, we compute SSIM and LPIPS between the input and front-view texture rendering. The best and the second best methods are highlighted in bold and underlined respectively. Note that * denotes models that have been trained on commercial 3D human scans while the others are trained on the public THuman2.0 dataset.
@@
Note that $\nabla_{\mathbf{x}}$ indicates numerical finite differences at point $\mathbf{x}$ and $\lambda_{n}$ is a hyperparameter.
During inference, we use the input image $I^{F}$ and the back-view image $\tilde{I}^{B}$ obtained from Sec. 3.1 to reconstruct 3D mesh and textures. First, we align both images with the estimated body mesh $\mathcal{M}$ to ensure that image features can be properly queried around the 3D anchor. We adopt a similar strategy with [54] to optimize the scale and the offset of the body mesh with silhouette and 2D joint errors. Finally, we perform the marching cube algorithm [29] by querying SDF and RGB values within a dense voxel grid via Eq. 3 (see Fig. 2 right).
@@ 4.1 Experimental Setup
Previous work [2, 44, 45] relied on training data from commercial datasets such as RenderPeople [1]. While these datasets offer high-quality textured meshes, they also limit reproducibility due to limited accessibility. For fair comparisons, we follow ICON [54] by training our method on the public 3D dataset THuman2.0 [56] and using the CAPE [30] dataset for evaluation. However, we observed potential biases in the evaluation due to the low-res ground-truth meshes and image rendering defects in the CAPE dataset (for a detailed discussion, please refer to Supp.). Consequently, we further create a new benchmark that evaluates the baselines on a higher-quality 3D human dataset CustomHumans [14]. In the following, we provide a summary of the datasets used in our experiments:
  • THuman2.0 [56] contains approximately 500 scans of humans wearing 150 different garments in various poses. We use these 3D scans to train our pipeline.
  • CAPE [30] contains 15 subjects in 8 types of tight outfits. The test set, provided by ICON, consists of 300 images rendered from three view angles. We use this dataset to conduct the 3D quantitative experiment (Sec. 4.2).
  • CustomHumans [14] contains more than 600 high-quality scans of 80 subjects in 120 different garments and varied poses. We randomly selected 60 subjects and rendered 240 test images from four view angles. We use this test data for all quantitative experiments, user studies, and ablation studies. (Sec. 4.2Sec. 4.4)
@@
We follow the evaluation protocol in OccNet [31] and ICON [54] to compute 3D metrics Chamfer distance (CD), normal consistency (NC), and f-Score [49] on the generated meshes. To evaluate mesh texture, we report structure similarity (SSIM[52], LPIPS [59]. In user studies, 30 participants rank the meshes obtained by four different methods. We report the average ranking ranging from 1 (best) to 4 (worst).

### v2

# SiTH: Single-view Textured Human Reconstruction with Image-Conditioned Diffusion

@@ Abstract
A long-standing goal of 3D human reconstruction is to create lifelike and fully detailed 3D humans from single-view images. The main challenge lies in inferring unknown body shapes, appearances, and clothing details in areas not visible in the images. To address this, we propose SiTH, a novel pipeline that uniquely integrates an image-conditioned diffusion model into a 3D mesh reconstruction workflow. At the core of our method lies the decomposition of the challenging single-view reconstruction problem into generative hallucination and reconstruction subproblems. For the former, we employ a powerful generative diffusion model to hallucinate unseen back-view appearance based on the input images. For the latter, we leverage skinned body meshes as guidance to recover full-body texture meshes from the input and back-view images.  SiTH requires as few as 500 3D human scans for training while maintaining its generality and robustness to diverse images. Extensive evaluations on two 3D human benchmarks, including our newly created one, highlighted our method’s superior accuracy and perceptual quality in 3D textured human reconstruction.
Figure 1: Single-view textured human reconstruction. SiTH is a novel pipeline for creating high-quality and fully textured 3D human meshes from single images. We first hallucinate back-view appearances through an image-conditioned diffusion model, followed by the reconstruction of full-body textured meshes using both the front and back-view images. Our pipeline enables the creation of lifelike and diverse 3D humans from unseen photos (left) and AI-generated images (right).
@@ 1 Introduction
Reconstructing a fully textured human mesh from a single-view image presents an ill-posed problem with two major challenges. Firstly, the appearance information required for generating texture in unobserved regions is missing. Secondly, 3D information for mesh reconstruction, such as depth, surface, and body pose, becomes ambiguous in a 2D image. Previous efforts [60, 79, 4] attempted to tackle these challenges in a data-driven manner, focusing on training neural networks with image-mesh pairs. However, these approaches struggle with images featuring unseen appearances or poses, due to limited 3D human training data. More recent studies [61, 79, 73] introduced additional 3D reasoning modules to enhance robustness against unseen poses. Yet, generating realistic and full-body textures from unseen appearances still remains an unsolved problem.
@@
To address the above challenges, we propose SiTH, a novel pipeline that integrates an image-conditioned diffusion model to reconstruct lifelike 3D textured humans from monocular images. At the core of our approach is the decomposition of the challenging single-view problem into two subproblems: generative back-view hallucination and mesh reconstruction. This decomposition enables us to exploit the generative capability of pretrained diffusion models to guide full-body mesh and texture reconstruction. The workflow is depicted in Fig. 1. Given a front-view image, the first stage involves hallucinating a perceptually consistent back-view image using image-conditioned diffusion. The second stage reconstructs full-body mesh and texture, utilizing both the front and back-view images as guidance.
@@
More specifically, we employ the generative capabilities of pretrained diffusion models (e.g. Stable Diffusion [57]) to infer unobserved back-view appearances for full-body 3D reconstruction. The primary challenge in ensuring the realism of 3D meshes lies in generating images that depict spatially aligned body shapes and perceptually consistent appearances with the input images. While diffusion models demonstrate impressive generative abilities with text conditioning, they are limited in producing desired back-view images using the frontal images as image conditions. To overcome this, we adapt the network architecture to enable conditioning on frontal images and introduce additional trainable components following ControlNet [77] to provide pose and mask control. To fully tailor this model to our task while retaining its original generative power, we carefully fine-tune the diffusion model using multi-view images rendered from 3D human scans. Complementing this generative model, we develop a mesh reconstruction module to recover full-body textured mesh from front and back-view images. We follow prior work in handling 3D ambiguity through normal [61] and skinned body [73, 79] guidance. It is worth noting that the models for both subproblems are trained using the same public THuman2.0 [75] dataset, which consists of as few as 500 scans.
To advance research in single-view human reconstruction, we created a new benchmark based on the high-quality CustomHumans [22] dataset and conducted comprehensive evaluations against state-of-the-art methods. Compared to existing end-to-end methods [60, 79, 4], our two-stage pipeline can recover full-body textured meshes, including back-view details, and demonstrates robustness to unseen images. In contrast to time-intensive diffusion-based optimization methods [41, 35, 24], our pipeline efficiently produces high-quality textured meshes in under two minutes. Moreover, we explored applications combining text-guided diffusion models, showing SiTH’s versatility in 3D human creation. Our contributions are summarized as follows:
  • We introduce SiTH, a single-view human reconstruction pipeline capable of producing high-quality, fully textured 3D human meshes within two minutes.
  • Through decomposing the single-view reconstruction task, SiTH can be efficiently trained with public 3D human scans and is more robust to unseen images.
  • We establish a new benchmark featuring more diverse subjects for evaluating textured human reconstruction.
@@ 2 Related Work
Reconstructing 3D humans from monocular inputs [67, 71, 26, 14, 17, 62, 28, 29] has gained more popularity in research. In this context, we focus on methods that recover 3D human shapes, garments, and textures from a single image. As a seminal work, Saito et al. [60] first proposed a data-driven method with pixel-aligned features and neural fields [72]. Its follow-up work PIFuHD [61] further improved this framework with high-res normal guidance. Later approaches extended this framework with additional human body priors. For instance, PaMIR [79] and ICON [73] utilized skinned body models [42, 51] to guide 3D reconstruction. ARCH [25], ARCH++ [21], and CAR [39] transformed global coordinates into the canonical coordinates to allow for reposing. PHOHRUM [4] and S3F [12] further disentangled shading and albedo to enable relighting. Another line of work replaced the neural representations with conventional Poisson surface reconstruction [34, 33]. ECON [74] and 2K2K [18] trained normal and depth predictors to generate front and back 2.5D point clouds. The human mesh is obtained by fusing these point clouds with body priors and 3D heuristics. However, none of these methods produce realistic full-body texture and geometry in the unobserved regions. Our pipeline addresses this problem by incorporating a generative diffusion model into the 3D human reconstruction workflow.
@@
Diffusion models [57, 59, 56, 52] trained with large collections of images have demonstrated unprecedented capability in creating 3D objects from text prompts. Most prior work [53, 68, 40, 47, 19, 11] followed an optimization workflow to update 3D representations (e.g. NeRF [48], SDF tetrahedron [63]) via neural rendering [66] and a score distillation sampling (SDS) [53] loss. While some methods [76, 10, 27, 64, 3, 24] applied this workflow to human bodies, they cannot produce accurate human bodies and appearances due to the ambiguity of text-conditioning. More recent work [41, 54] also tried to extend this workflow with more accurate image-conditioning. However, we show that they struggle to recover human clothing details and require a long optimization time. Most related to our work is Chupa [35], which also decomposes its pipeline into two stages. Note that Chupa is an optimization-based approach that relies on texts and cannot model colors. We address these issues by introducing an image-conditioning strategy and model. Most importantly, our method swiftly reconstructs full-texture human meshes without any optimization process.
@@
Foundation models [13, 8, 20, 36] trained on large-scale datasets have been shown to be adaptable to various downstream tasks. Following this trend, pretrained diffusion models [57, 59, 56, 52] have become common backbones for generative modeling. For instance, they can be customized by finetuning with a small collection of images [23, 37, 58]. ControlNet [77] introduced additional trainable plugins to enable image conditioning such as body skeletons. While these strategies have been widely adopted, none of them directly fit our objective. More relevant to our task is DreamPose [31], which utilizes DensePose [16] images as conditions to repose input images. However, it cannot handle out-of-distribution images due to overfitting. Similarly, Zero-1-to-3 [41] finetunes a diffusion model with multi-view images to allow for viewpoint control. However, we show that viewpoint conditioning is not sufficient for generating consistent human bodies. Our model addresses this issue by providing accurate body pose and mask conditions for back-view hallucination.
@@ 3.1 Back-view Hallucination
Simply applying the LDM architecture to our task is not sufficient since our goal is to learn a conditional distribution of back-view images given an input conditional image. To this end, we make several adaptations to allow for image-conditioning as shown in Fig. 3. First, we utilize the pretrained CLIP [55] image encoder and VAE encoder $\mathcal{E}$ to extract image features from the front-view image (i.e., $I^{F}$). These image features are used for conditioning the LDM, ensuring the output image shares a consistent appearance with the input image. Second, we follow the idea of ControlNet [77] and propose to use a UV map ($I^{B}_{UV}\in\mathbb{R}^{H\times W\times 3}$) and a silhouette mask ($I^{B}_{M}\in\mathbb{R}^{H\times W}$) from the back view as additional conditions. These conditional signals provide additional information that ensures the output image has a similar body shape and pose to the conditional input image.
@@
Another challenge in training an image-conditioned LDM is data. Training the model from scratch is infeasible due to the requirement of a large number of paired images rendered from 3D textured human scans. Inspired by the concept of learning from large-scale pretraining [13, 20], we build our image-conditioned LDM on top of a pretrained diffusion U-Net [57]. We utilize the finetuning strategy [37, 77] to optimize cross-attention layers and ControlNet parameters while keeping most of the other parameters frozen (see Fig. 3). The design and training strategy of our image-conditioned diffusion model enables hallucinating plausible back-view images that are cosistent with the frontal inputs.
@@ 3.2 Human Mesh Reconstruction
$\mathbf{f}_{r,x}$ $=\mathcal{B}(\mathbf{f}_{r},\pi(\mathbf{x}))=\mathcal{B}(G_{r}(I),\pi(\mathbf{x})),$
@@
$\mathbf{r}_{x}$ $=H_{r}(\mathbf{f}_{r,x}^{F},\mathbf{f}_{r,x}^{B},\mathbf{p}).$
CAPE [45] CustomHuman [22]
Method
CD: P-to-S /
S-to-P (cm)$\downarrow$
@@
LPIPS: F
(×$(\times 10^{-2})\downarrow$
CD: P-to-S /
S-to-P (cm)$\downarrow$
[…]
LPIPS: F / B
(×$(\times 10^{-2})\downarrow$
PIFu [60] 2.368 / 3.763 0.778 33.842 2.720 2.209 / 2.582 0.805 34.881 6.073 / 8.496
PIFuHD [61] 2.401 / 3.522 0.772 35.706 - 2.107 / 2.228 0.804 39.076 -
PaMIR [79] 2.190 / 2.806 0.804 36.725 2.085 2.181 / 2.507 0.813 35.847 4.646 / 7.152
2K2K [18] 2.478 / 3.683 0.782 28.700 - 2.488 / 3.292 0.796 30.186 -
FOF [15] 2.196 / 4.040 0.777 34.227 - 2.079 / 2.644 0.808 36.013 -
ICON [73] 2.516 / 3.079 0.786 29.630 - 2.256 / 2.795 0.791 30.437 -
ECON [74] 2.475 / 2.970 0.788 30.488 - 2.483 / 2.680 0.797 30.894 -
SiTH (Ours) 1.899 / 2.261 0.816 37.763 1.977 1.871 / 2.045 0.826 37.029 3.929 / 6.803
Table 1: Single-view human reconstruction benchmarks. We report Chamfer distance (CD), normal consistency (NC), and f-score between ground truth and predicted meshes. To evaluate texture reconstruction quality, we compute LPIPS between the image rendering of GT and generated textures. The best and the second best methods are highlighted in bold and underlined respectively. Note that gray color denotes models trained on more commercial 3D human scans while the others are trained on the public THuman2.0 dataset.
@@
Note that $\nabla_{\mathbf{x}}$ indicates numerical finite differences for computing local normals at point $\mathbf{x}$ and $\lambda_{n}$ is a hyperparameter.
During inference, we use the input image $I^{F}$ and the back-view image $\tilde{I}^{B}$ obtained from Sec. 3.1 to reconstruct 3D mesh and textures. First, we align both images with the estimated body mesh $\mathcal{M}$ to ensure that image features can be properly queried around the 3D anchor. We adopt a similar strategy of SMPLify [7] to optimize the scale and the offset of the body mesh with silhouette and 2D joint errors. Finally, we perform the marching cube algorithm [43] by querying SDF and RGB values within a dense voxel grid via Eq. 3 (see Fig. 2 right).
@@ 4.1 Experimental Setup
Previous work relied on training data from commercial datasets such as RenderPeople [1]. While these datasets offer high-quality textured meshes, they also limit reproducibility due to limited accessibility. For fair comparisons, we follow ICON [73] by training our method on the public 3D dataset THuman2.0 [75] and using the CAPE [45] dataset for evaluation. However, we observed potential biases in the evaluation due to the low-res ground-truth meshes and image rendering defects in the CAPE dataset (for a detailed discussion, please refer to Supp-Sec. 6). Consequently, we further create a new benchmark that evaluates the baselines on a higher-quality 3D human dataset CustomHumans [22]. In the following, we provide a summary of the datasets used in our experiments:
  • THuman2.0 [75] contains approximately 500 scans of humans wearing 150 different garments in various poses. We use these 3D scans as the training data.
  • CAPE [45] contains 15 subjects in 8 types of tight outfits. The test set, provided by ICON, consists of 100 meshes. We use CAPE for the quantitative evaluation (Sec. 4.2).
  • CustomHumans [22] contains 600 higher-quality scans of 80 subjects in 120 different garments and varied poses. We selected 60 subjects for all quantitative experiments, user studies, and ablation studies. (Sec. 4.2Sec. 4.4)
@@
We follow the evaluation protocol in OccNet [46] and ICON [73] to compute 3D metrics Chamfer distance (CD), normal consistency (NC), and f-Score [65] on the generated meshes. To evaluate reconstructed mesh texture, we report LPIPS [78] of front and back texture rendering. In user studies, 30 participants rank the meshes obtained by four different methods. We report the average ranking ranging from 1 (best) to 4 (worst).